# The Common Data Format (CDF): A Standardized Format for Match-Data in Football (Soccer)

### v3

@@ Appendix C Position Labels
Formations as well as player positions are subjective. Multiple approaches have tried to objectify the discussions around formations and positions using tracking data [Bialkowski2014, Bialkowski2015, Bauer2021_Formations_IEEE]. The take away learnings from these studies is that playing positions should be considered dynamic, and thus are constantly changing within a match. However, static formations defined for at least part of the matches are a relevant simplification used by experts. In order to create a comparability among vendors, we map player positions from different sources to standardized positions. Note that these positions describe pre-match assigned positions.

### v4

@@ 6 Conclusions
Acknowledgments
JD received support from the Flemish Government under the “Onderzoeksprogramma Artifciële Intelligentie (AI) Vlaanderen” program.
@@ Appendix C Position Labels
Formations as well as player positions are subjective. Multiple approaches have tried to objectify the discussions around formations and positions using tracking data [8, 7, 4]. The take away learnings from these studies is that playing positions should be considered dynamic, and thus are constantly changing within a match. However, static formations defined for at least part of the matches are a relevant simplification used by experts. In order to create a comparability among vendors, we map player positions from different sources to standardized positions. Note that these positions describe pre-match assigned positions.